OpenStudy (anonymous):

Which of the following is a solution of y - x < -3? (6, 2) (2, 6) (2, -1)

jimthompson5910 (jim_thompson5910):

If (6,2) was a solution to y - x < -3, then x = 6 and y = 2 should make y - x < -3 true

jimthompson5910 (jim_thompson5910):

y - x < -3 2 - 6 < -3 ... plug in (x,y) = (6,2) -4 < -3 ... is this true?